A CIDR block extends from 204.206.160.0 to 204.206.175.255.

a. How many bits are in the net id?

b. How many bits are in the host id?

c. What is the mask in slash and binary notation?

d. What is the maximum number of valid IP addresses that are available?

e. If I want 19 subnets how does the subnet mask change? Show the first 2 in slash notation.

f.What is the maximum number of valid IP addresses available per subnet?
Answer #1

Total number of addresses are : 16*256= 2^12

a. Bits in network id = 12 bits

b. bits in host id= 32-n bits = 20 bits

c. Mask is used to determine how an IP (Internet Protocol) message sent by an IP host (such as your PC) will be routed

subnet mask : 255.240.0.0

we can represent as 204.206.160.0/12

1111111111110000000000000000000

d. Total addresses available are 1,048,576

Valid addresses are 1,048,574

e. Subnet mask would be /17

or 255.255.128.0

f. number of valid IP addresses :32,766